OpenStudy (aravindg):

guys giv me general sormula sin 2x=cos 2x

OpenStudy (anonymous):

finally one we can do!

OpenStudy (aravindg):

lol

OpenStudy (anonymous):

we know \[2x=\frac{\pi}{4}\] or \[2x=\frac{5\pi}{4}\] etc that is \[2x=\frac{(2k+1)\pi}{4}\]

OpenStudy (anonymous):

solve for x in one step and be happy we can actually get an answer

OpenStudy (aravindg):

???

OpenStudy (aravindg):

i need general solution between 0<x<3\[\Pi\]/4

OpenStudy (anonymous):

\[x=\frac{(2k+1)\pi}{8},x\in \mathbb Z\]

OpenStudy (anonymous):

ooh i finally got the "zahl" right!

OpenStudy (aravindg):

0<x<3Π/4

OpenStudy (aravindg):

ur answer is wrong